OpenStudy (anonymous):

Solve S = 4v2 for v.

jimthompson5910 (jim_thompson5910):

I'm assuming the equation is \[\large S = 4v^2\]

OpenStudy (anonymous):

yes i was a little confused becasue this is all they gave

OpenStudy (anonymous):

you want the equation in the form of v= _____

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

First divide both sides by 4 to get \[\large S = 4v^2\] \[\large \frac{S}{4} = v^2\] \[\large v^2 = \frac{S}{4}\] what's your next step?

OpenStudy (anonymous):

next step be square rooting it? im really lost in this whole lesson

jimthompson5910 (jim_thompson5910):

good, that undoes the square exponent

jimthompson5910 (jim_thompson5910):

keep in mind that if you have something like x^2 = 4 then x = 2 or x = -2 since both are solutions (when you replace x with either value)

jimthompson5910 (jim_thompson5910):

so this means \[\large v^2 = \frac{S}{4}\] turns into \[\large v = \pm\sqrt{\frac{S}{4}}\] after you take the square root of both sides

jimthompson5910 (jim_thompson5910):

you can optionally write \[\large v = \pm\sqrt{\frac{S}{4}}\] as \[\large v = \sqrt{\frac{S}{4}} \ \text{ or } \ v = -\sqrt{\frac{S}{4}} \] but they mean the same thing

jimthompson5910 (jim_thompson5910):

the first way is more compact

OpenStudy (anonymous):

so it can be either negative or positively written?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ok wow thank u! so much!

jimthompson5910 (jim_thompson5910):

because (2)^2 = 4 and (-2)^2 = 4 so this shows you that x^2 = 4 has two solutions x = -2 or x = 2, which is really x = plus/minus 2

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Solve M = 2x + 3y for y I also have a question like this, think you could help? explaining it helps I do online classes and really do not have any well explaining

jimthompson5910 (jim_thompson5910):

how do you undo the +2x

OpenStudy (anonymous):

wudnt u have to cancel it out

jimthompson5910 (jim_thompson5910):

how would you cancel out 2x on the right side

OpenStudy (anonymous):

subtract 2x from the y side also. Im not really sure

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\large M = 2x + 3y\] \[\large M - 2x = 3y\] \[\large 3y = M - 2x\] what's next?

OpenStudy (anonymous):

would we add 2, to the 2x to cancel it out and get the y all by itself

jimthompson5910 (jim_thompson5910):

3y means 3 times y we move that 3 over by undoing the "times 3" what will undo the "times 3"?

OpenStudy (anonymous):

wud we do 3x2 to find out what we re multiplying? im not sure

OpenStudy (anonymous):

i never no which numbers to fill in to get the rest of the problem

jimthompson5910 (jim_thompson5910):

let's say we start with the number 5 then we multiply it by 3 to get 5*3 = 15 how do we undo this and go back to 5?

OpenStudy (anonymous):

15-10 wud get us back to 5

jimthompson5910 (jim_thompson5910):

true, but what operation undoes multiplication what is the opposite of multiplication?

OpenStudy (anonymous):

division

jimthompson5910 (jim_thompson5910):

so if we multiply by 3 to do the complete opposite, we divide by 3

OpenStudy (anonymous):

15 divided by 3

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

this means we divide both sides by 3

jimthompson5910 (jim_thompson5910):

so the complete step by step picture looks like this \[\large M = 2x + 3y\] \[\large M - 2x = 3y\] \[\large 3y = M - 2x\] \[\large y = \frac{M - 2x}{3}\]

jimthompson5910 (jim_thompson5910):

so at first, we undid the addition of 2x...so we subtracted 2x from both sides then we undid the multiplication of 3...which means we divided both sides by 3

OpenStudy (anonymous):

we subtracted 2x fromthe 2x and 3y then divided 33 and divided it from 2x?

jimthompson5910 (jim_thompson5910):

yes we go from 2x+3y to just 3y when you subtract off 2x you must do the same to the other side, which explains how it goes from M to M - 2x

jimthompson5910 (jim_thompson5910):

we now have \[\large 3y = M - 2x\] then you divide both sides by 3 so 3y divided by 3 = y since the 3s cancel and dividing the right side by 3 means you can leave it as \[\large y = \frac{M - 2x}{3}\] since not much simplification is done anyway

OpenStudy (anonymous):

so divide 2x and the 3 thts how we get it

jimthompson5910 (jim_thompson5910):

correct and you leave the right side as it is

OpenStudy (anonymous):

so this wud be everything after simplifying Y=M-2 over 3

jimthompson5910 (jim_thompson5910):

yep \[\large y = \frac{M - 2x}{3}\]

jimthompson5910 (jim_thompson5910):

oh it should be -2x and not just -2

OpenStudy (anonymous):

2x ok got it thank u

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

:)